# A hypersphere-like non-Abelian Yang monopole and its topological characterization

### v1

@@ Abstract
where $|\psi_{-}^{\alpha,\beta}\rangle$ are two eigenstates corresponding to the lower eivenvalues and $\langle\tilde{\psi}_{-}^{\alpha,\beta}|$ are the normalized left eigenstate of $|\psi_{-}^{\alpha,\beta}\rangle$. Without loss of generality, we set $\kappa=1$ and assume the parameter of $\vec{q}=R\{\textrm{sin}\theta_{1}\textrm{cos}\theta_{2}\textrm{cos}\phi_{1},\textrm{sin}\theta_{1}\textrm{cos}\theta_{2}\textrm{sin}\phi_{1},\textrm{sin}\theta_{1}\textrm{sin}\theta_{2}\textrm{cos}\phi_{2},\textrm{sin}\theta_{1}\\ \textrm{sin}\theta_{2}\textrm{sin}\phi_{2},\textrm{cos}\theta_{1}\}$ to construct a manifold with the four angular coordinates $\{\theta_{1},\theta_{2},\phi_{1},\phi_{2}\}$ and the radial coordinate $R$. $C_{2}$ is calculated according to Eqs. (2) and (3) (see Supplemental Material), whose values with respect to $R$ is shown in Fig. 2. $C_{2}$ is equal to 1 or 0, dependent on whether the 5D parameter manifold encloses ($R>1$) or does not enclose ($R<1$) the EHS, with the critical boundary at $R=1$ where the $C_{2}$ jumps between the two values, revealing topological transition inherent in such a 4D NH model. Note that the topological charge here is entirely carried by the whole EHS.
@@
Figure 3: The characterization of the Wilson loop and of its relations. (a) Parameter manifold (projected onto yellow sphere) enclosing the EHS (projected onto orange sphere). The traversing path $\mathcal{L}$ to acquire (b) the Wilson loop $W_{\mathcal{L}}$ at distinct phase angle $\theta_{2}$. (c) Trajectory of loop $\mathcal{L}$ when the parameter manifold is within the EHS and (d) its corresponding $W_{\mathcal{L}}$ versus $\theta_{2}$. Expectation values of the three Pauli operators $\langle\sigma_{j}\rangle$ (left) and of their square summation $\langle\sigma^{2}\rangle$ (right) under the system evolution with respect to angle $\phi_{1}$ (e) for the traversing path $\mathcal{L}$ with $\theta_{2}=\pi/4$. (f) Minimal values of $W_{\mathcal{L}}$ versus the radius $R$ of the parameter manifold. (g) Real and (h) imaginary parts of the spectra on the Riemann surface for path $\mathcal{L}$ encircling the ring which is the projection of the EHS in $\{q_{1},q_{2},q_{4}\}$ subspace. (i) Both real and imaginary parts of the spectra, where the EP is the projection of the EHS in $q_{1,2}$ axis, and (j) the corresponding $W_{\mathcal{L}}$ as function of $\Delta/R$, respectively, when path $\mathcal{L}$ encircling ($\Delta/R<2$) and non-encircling ($\Delta/R>2$) the EHS.
From other point of view, we can also employ the Wilczek-Zee (WZ) geometric phase to characterize the exceptional physics of such an EHS. The unitary evolution operator characterizing the system dynamics can be described by the WZ phase factor [16]
@@
where $\mathcal{P}$ denotes the path-ordered integral along the path $\mathcal{L}$. The Wilson loop can be obtained by tracing out the $U_{\mathcal{L}}$: $W_{\mathcal{L}}=\textrm{tr}(U_{\mathcal{L}})$. For the model in Eq. (1), we parameterize the space as $\vec{q}=R\{\textrm{cos}\theta_{2}\textrm{cos}\phi_{1},\textrm{cos}\theta_{2}\textrm{sin}\phi_{1},\textrm{sin}\theta_{2},0,0\}$ ($\phi_{1}\in[0,2\pi]$). When $R>1$, the closed loop $\mathcal{L}$ traced by the control vector $\vec{q}$ is depicted in Fig. 3(a). Fig. 3(b) shows the Wilson loop, defined as $W_{\mathcal{L}}=\textrm{tr}\left[\mathcal{P}\textrm{exp}\left(i\int_{0}^{2\pi}A_{\phi_{1}}\textrm{d}{\phi_{1}}\right)\right]$, versus $\theta_{2}(\in[0,\pi])$. At $\theta_{2}=0$, $\int A^{11}=-\int A^{22}=\pi$, which corresponds to the global minimum $W_{\mathcal{L}}=-2$; while at $\theta_{2}=\pi/2$, $\int A^{11}=\int A^{22}=0$, corresponding to the global maximum $W_{\mathcal{L}}=2$ (see Supplemental Material). When $R<1$, the parameter manifold does not enclose the EHS, the trajectory of the closed loop $\mathcal{L}$} and the corresponding $W_{\mathcal{L}}$ are depicted in Figs. 3(c) and 3(d), respectively. Fig. 3(e) (left) displays the evolution of the expectation values of Pauli operators $\langle\sigma_{j}\rangle$ ($j=x,y,z$) within the degenerate ground subspace as the parameter $\phi_{1}$ traverses a closed loop from 0 to $2\pi$, with the system starting from one of the degenerate ground states at $\theta_{2}=\pi/4$. This leads to their square summation $\langle\sigma^{2}\rangle\equiv\langle\sum\sigma_{j}^{2}\rangle<1$, as shown in Fig. 3(e) (right). This demonstrates the non-Abelian feature of our structured 4D NH system, in stark contrast with its Abelian counterpart of $\langle\sigma^{2}\rangle=1$. Fig. 3(f) shows the minimal $W_{\mathcal{L}}$ versus the radius R of the parameter manifold over $\theta_{2}\in[0,\pi]$. Remarkably, it is found that, the minimal $W_{\mathcal{L}}$ remains 2-2 and keeps stable when the parameter manifold encloses the EHS; while it goes through a sharp increase when the parameter manifold crosses the EHS, and then smoothly increases, tending to $2$. It should be noticed that the Wilson loop $W_{\mathcal{L}}$ is not a topological quantity, and thus a sharp change in its values does not mean that the system undergoes a topological transition, which only represents a manifestation of the inherent geometric feature.

### v2

@@ Abstract
where $|\psi_{-}^{\alpha,\beta}\rangle$ are two eigenstates corresponding to the lower eivenvalues and $\langle\tilde{\psi}_{-}^{\alpha,\beta}|$ are the normalized left eigenstate of $|\psi_{-}^{\alpha,\beta}\rangle$. Without loss of generality, we set $\kappa=1$ and assume the parameter of $\vec{q}=R\{\textrm{sin}\theta_{1}\textrm{sin}\theta_{2}\textrm{cos}\phi_{2},\textrm{sin}\theta_{1}\textrm{cos}\theta_{2}\textrm{cos}\phi_{1},\textrm{sin}\theta_{1}\textrm{cos}\theta_{2}\textrm{sin}\phi_{1},\textrm{cos}\theta_{1},\\ \textrm{sin}\theta_{1}\textrm{sin}\theta_{2}\textrm{sin}\phi_{2}\}$ to construct a manifold with the four angular coordinates $\{\theta_{1},\theta_{2},\phi_{1},\phi_{2}\}$ and the radial coordinate $R$. $C_{2}$ is calculated according to Eqs. (2) and (3) (see Supplemental Material), whose values with respect to $R$ is shown in Fig. 2. $C_{2}$ is equal to 1 or 0, dependent on whether the 5D parameter manifold encloses ($R>1$) or does not enclose ($R<1$) the EHS, with the critical boundary at $R=1$ where the $C_{2}$ jumps between the two values, revealing topological transition inherent in such a 4D NH model. Note that the topological charge here is entirely carried by the whole EHS.
@@
Figure 3: The characterization of the Wilson loop and of its relations. (a) Parameter manifold (projected onto yellow sphere) enclosing the EHS (projected onto orange sphere). The traversing path $\mathcal{L}$ to acquire (b) the Wilson loop $W_{\mathcal{L}}$ at distinct phase angle $\theta_{2}$. (c) Trajectory of loop $\mathcal{L}$ when the parameter manifold is within the EHS and (d) its corresponding $W_{\mathcal{L}}$ versus $\theta_{2}$. Expectation values of the three Pauli operators $\langle\sigma_{j}\rangle$ (left) and of their square summation $\langle\sigma^{2}\rangle$ (right) under the system evolution with respect to angle $\phi_{1}$ (e) for the traversing path $\mathcal{L}$ with $\theta_{2}=\pi/4$. (f) Minimal values of $W_{\mathcal{L}}$ versus the radius $R$ of the parameter manifold. (g) Real and (h) imaginary parts of the spectra on the Riemann surface for path $\mathcal{L}$ encircling the ring which is the projection of the EHS in $\{q_{1},q_{2},q_{4}\}$ subspace. (i) Both real and imaginary parts of the spectra, where the EP is the projection of the EHS in $q_{1,2}$ axis, and (j) the corresponding $W_{\mathcal{L}}$ as function of $\Delta/R$, respectively, when path $\mathcal{L}$ encircling ($\Delta/R<2$) and non-encircling ($\Delta/R>2$) the EHS. The shadow region marks the critical boundary at $\Delta/R=2$ where the transition happens.
From other point of view, we can also employ the WZ geometric phase to characterize the exceptional physics of such an EHS. The unitary evolution operator characterizing the system dynamics can be described by the WZ phase factor [16]
@@
where $\mathcal{P}$ denotes the path-ordered integral along the path $\mathcal{L}$. The Wilson loop can be obtained by tracing out the $U_{\mathcal{L}}$: $W_{\mathcal{L}}=\textrm{tr}(U_{\mathcal{L}})$. For the model in Eq. (1), we parameterize the space as $\vec{q}=R\{\textrm{sin}\theta_{2},\textrm{cos}\theta_{2}\textrm{cos}\phi_{1},\textrm{cos}\theta_{2}\textrm{sin}\phi_{1},0,0\}$ ($\phi_{1}\in[0,2\pi]$). When $R>1$, the closed loop $\mathcal{L}$ traced by the control vector $\vec{q}$ is depicted in Fig. 3(a). Fig. 3(b) shows the Wilson loop, defined as $W_{\mathcal{L}}=\textrm{tr}\left[\mathcal{P}\textrm{exp}\left(i\int_{0}^{2\pi}A_{\phi_{1}}\textrm{d}{\phi_{1}}\right)\right]$, versus $\theta_{2}(\in[0,\pi])$. At $\theta_{2}=0$, $\int A^{11}=-\int A^{22}=\pi$, which corresponds to the global minimum $W_{\mathcal{L}}=-2$; while at $\theta_{2}=\pi/2$, $\int A^{11}=\int A^{22}=0$, corresponding to the global maximum $W_{\mathcal{L}}=2$ (see Supplemental Material). When $R<1$, the parameter manifold does not enclose the EHS, the trajectory of the closed loop $\mathcal{L}$} and the corresponding $W_{\mathcal{L}}$ are depicted in Figs. 3(c) and 3(d), respectively. Fig. 3(e) (left) displays the evolution of the expectation values of Pauli operators $\langle\sigma_{j}\rangle$ ($j=x,y,z$) within the degenerate ground subspace as the parameter $\phi_{1}$ traverses a closed loop from 0 to $2\pi$, with the system starting from one of the degenerate ground states at $\theta_{2}=\pi/4$. This leads to their square summation $\langle\sigma^{2}\rangle\equiv\langle\sum\sigma_{j}^{2}\rangle<1$, as shown in Fig. 3(e) (right). This demonstrates the non-Abelian feature of our structured 4D NH system, in stark contrast with its Abelian counterpart of $\langle\sigma^{2}\rangle=1$. Fig. 3(f) shows the minimal $W_{\mathcal{L}}$ versus the radius $R$ of the parameter manifold over $\theta_{2}\in[0,\pi]$. Remarkably, it is found that, the minimal $W_{\mathcal{L}}$ remains 2-2 and keeps stable when the parameter manifold encloses the EHS; while it goes through a sharp increase when the parameter manifold crosses the EHS, and then smoothly increases, tending to $2$. It should be noticed that the Wilson loop $W_{\mathcal{L}}$ is not a topological quantity, and thus a sharp change in its values does not mean that the system undergoes a topological transition, which only represents a manifestation of the inherent geometric feature.